5. Suppose that you and your dorm mates decide to construct a Halloween Party Price Index. After careful research you have come up with the following information:

Halloween Party Price Index

item

quantity

price in 2018

price in 2019

36 oz. bag of candy bars

4

$2.00

$4.75

gallon of apple cider

2

$2.50

$3.00

one-half keg of beer

1

$90.00

$100.00

Teenage Mutant Ninja Turtles costumes

4

$14.00

$18.00

16 oz. plastic cups

100

$0.03

$0.05

13 oz. bag Tostitos

6

$3.00

$4.00

24 oz. jar salsa

2

$4.00

$3.00

pumpkins

4

$2.00

$2.50

candles

4

$1.00

$2.00

Assume that 2018 is the base year.

a. What is the cost of a Halloween Party in 2018?

b. What is the cost of a Halloween Party in 2019?

c. Construct a Halloween Party Price Index (HPPI) for 2018.

d. Construct a Halloween Party Price Index (HPPI) for 2019.

e. Using the HPPI, what is the rate of inflation?
Answer #1

a. The cost of a Halloween Party in 2018 = Sum of (Quantity*Respective prices in 2018) = (4*2) + (2*2.5) + (1*90) + (4*14) +
(100*0.03) + (6*3) + (2*4) + (4*2) + (4*1) = 8 + 5 + 90 + 56 + 3 + 18 + 8 + 8 + 4 = 200

b. The cost of a Halloween Party in 2019 = Sum of (Quantity*Respective prices in 2019) = (4*4.75) + (2*3) + (1*100) + (4*18) +
(100*0.05) + (6*4) + (2*3) + (4*2.5) + (4*2) = 19 + 6 + 100 + 72 + 5 + 24 + 6 + 10 + 8 = 250

c. Halloween Party Price Index (HPPI) for 2018 = (Cost in 2018/Cost in base year)*100 = (200/200)*100 = 100

d. Halloween Party Price Index (HPPI) for 2019 = (Cost in 2019/Cost in base year)*100 = (250/200)*100 = 125

e. Inflation rate = [(HPPI for 2019-HPPI for 2018)/HPPI for 2018]*100 = [(125-100)/100]*100 = 25%